Charles must make a loan payment of $2,150 at the end of each quarter. His nominal annual loan rate is 2%, compounded quarterly. The original term of the loan was 13 years. How much did Charles originally borrow?
Answer #1

i =2%/4 = 0.5% per quarter

n = 13*4 = 52 quarters

Charles originally boorowed = 2150*(P/A,0.5%,52)

= 2150*45.68975

= 98233